OpenStudy (gabylovesyou):

What is the solution to the equation 14 x + 3 = 21? x ≈ −2.133 x ≈ −1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (gabylovesyou):

@jim_thompson5910

satellite73 (satellite73):

let me guess \[14^{x+3}=21\] right?

OpenStudy (gabylovesyou):

yes @satellite73 sorry cx

satellite73 (satellite73):

\[14^{x+3}=21\iff x+3=\frac{\log(21)}{\log(14)}\]

satellite73 (satellite73):

that is the change of base formula you still have to solve for \(x\) so \[x=\frac{\log(21)}{\log(14)}-3\] and a calculator

OpenStudy (gabylovesyou):

so the answer is B .. Thank you.

satellite73 (satellite73):

want me to check?

satellite73 (satellite73):

yeah you got it B

OpenStudy (gabylovesyou):

:)